OpenStudy (anonymous):

Solve for x. logx+log(x-15)=2

hartnn (hartnn):

hello :) what did you try?? tried combining those 2 logs?

OpenStudy (anonymous):

I know what you have to do, however, I am not sure how to do it :(

OpenStudy (anonymous):

I don't know how to combine the logs

hartnn (hartnn):

\(\log A + \log B = \log AB \) so how about log x + log (x-15) ?

OpenStudy (anonymous):

is it 20?

hartnn (hartnn):

you're asking if x = 20 ? how did you do all the steps so fast! :O

OpenStudy (anonymous):

I've been working on this problem for a really long time and I wasn't sure if I was doing the steps right so I posted on here. I think the answer is 20...I don't know what else it could be :(

hartnn (hartnn):

oh great! you've tried some steps... can you post them here, we'll correct then if required.

OpenStudy (anonymous):

Oh boy... I don't know where to begin! I'll write the most recent steps I've tried...one moment

OpenStudy (anonymous):

I used the formula \[a=\log _{b}(b ^{a})\]

OpenStudy (anonymous):

\[2= \log _{10}(10^2)\]

hartnn (hartnn):

for the right side? so, \(2 = \log 10^2 \)

hartnn (hartnn):

but for the left side we will need the formula I posted :)

hartnn (hartnn):

\(\log x + \log(x-15) = \log x\times(x-15) \\= \log 10^2\)

OpenStudy (anonymous):

\[\log _{10}(x)+\log _{10}(x-15)=\log _{10}(10^2)\]

hartnn (hartnn):

good

OpenStudy (anonymous):

\[\log _{10}(x(x-15))=\log _{10}(10^2)\]

hartnn (hartnn):

you get how \(\Large \log x + \log (x-15) = \log [x\times (x-15)]\) basically just a log property

OpenStudy (anonymous):

and the answeres are x=20 and x=-5 but we cant have a negative log so it must be 20

hartnn (hartnn):

if log A = log B, then A = B what that says is that you can just cancel out log on both sides! so we have x(x-15) = 10^2 got this?

OpenStudy (anonymous):

yes I just skipped the last steps at the end because they were really small and by then I already figured out my answers :)

hartnn (hartnn):

oh you solved it completely. excellent! x =20 is correct :)

OpenStudy (anonymous):

Thanks @hartnn , I appreciate your time :)

hartnn (hartnn):

welcome ^_^ happy to help :)